X Sorry, thats not correct. First find the present equivalent of Wills payments on June by discounting the September payment for three months to bring it back to June and adding it to the June payment. Then find the present equivalent on April by discounting the present equivalent on June for two months and adding it to the April payment. Finally, in the same way calculate the present equivalent of all four payments on January 15 Recall the formula for the (P/F, /%, N) factor: on! OKWill must make quarterly estimated income tax payments to the Internal Revenue Service. The amounts and timing of these payments are shown below. Also shown are the monthly interest rates that apply to each interval of time. What is the P-equivalent of Wills payments on January 15? Jan. 15 April 15 June 15 Sept. 15 mo mo mo $9,000 $11,000 $12,000 $9,000 The P-equivalent of Wills payments on January 15 is $40257 (Round to the nearest dollar.)This answer is incorrect, as well as 40912.

Please solve correctly, second time posting getting the incorrect answer.
Answer #1

Present value of Payment made on Jan15=Po=$9000

Future value of Payment made on April15=F1=$11000

Interest Rate for the Payment made on April15=i1=0.25%/month

Period from Jan 15=n1=3 months

We know that (P/F,i,N)=1/(1+i)n

(P/F,0.25%,3)=1/(1+0.25%)3=0.992537

P-Equivalent for payment on April 15=P1 =11000(P/F,0.25%,3)=11000*0.992537=$10917.91

Future value of Payment made on June 15=F2=$12000

Interest Rate for the Payment made on June15(from 15th april)=i2=0.50%/month

Period from April 15 to this payment=n2=2 months

(P/F,0.50%,2)(P/F,0.25%,3) =(1/(1+0.50%)2) (1/(1+0.25%)3)=0.990075*0.992537=0.982686

P-Equivalent for payment on June 15=P2= 12000(P/F,0.50%,2)(P/F,0.25%,3)

=12000*0.982686=11792.23

Future value of Payment made on Sept 15=F3=$9000

Interest Rate for the Payment made on Sept 15 (from 15th June)=i3=0.75%/month

Period from June 15 to this payment=n3=3 months

(P/F,0.75%,3) (P/F,0.50%,2)(P/F,0.25%,3)=(1/(1+0.75%)3)(1/(1+0.50%)2) (1/(1+0.25%)3)= 0.977833*0.982686=0.960903

P-Equivalent for payment on Sept 15=P3= (P/F,0.75%,3) (P/F,0.50%,2)(P/F,0.25%,3) =9000*0.960903=$8648.127

P-Equivalent=Po+P1+P2+P3=9000+10917.91+11792.23+8648.13=$40358.27